OpenStudy (anonymous):

Check all polynomials that are perfect square trinomials: 49x2 − 8x + 16 4a2 − 20a + 25 25b2 − 5b + 10 16x2 − 8x + 1

OpenStudy (nurali):

If ax² + bx + c is a perfect square, then the discriminant is 0. The discriminant = b² - 4ac 49x² − 8x + 16: discriminant = (-8)² - 4(49)(16) = -3072 4a² − 20a + 25: discriminant = (-20)² - 4(4)(25) = 0 25b² − 5b + 10: discriminant = (-5)² - 4(25)(10) = -975 16x² − 8x + 1: discriminant = (-8)² - 4(16)(1) = 0

OpenStudy (blurbendy):

The second and last ones